The area of a rectangle wall of a barn is 189 sq feet. Its length is 12 feet longer than the width. Find the length and width of the wall of the barn.
Thank you.

Answer by checkley77(12844)

L=W+12
189=LW
189=(W+12)W
189=W^2+12W
W^2+12W-189=0
(W-9)(W+21)=0
W-9=0
W=9 ANS. FOR THE WIDTH
L=9+12=21 ANS. FOR THE LENGTH.
PROOF:
189=9*21
189=189